# Leveraging Skills from Unlabeled Prior Data for Efficient Online Exploration

### v2

@@ Appendix A Compute Resources
Thus, in total the results in this paper required approximately $16,601$ GPU hours, or about 1.9 GPU years. Note this is an approximate upper bound, since not all methods required training checkpoints, and checkpoints were shared between different baselines that both uses trajectory skills or both used HILP skills.
@@ Appendix B VAE Architecture and hyperparameters
GRU Hidden Size 256 (AntMaze, Kitchen, VisualAntMaze)
512 (AntSoccer, HumanoidMaze, Scene, Cube)
GRU Layers 2 hidden layers (AntMaze, Kitchen, VisualAntMaze)
3 hidden layers (AntSoccer, HumanoidMaze, Scene, Cube)
KL Coefficient ($\beta)$ 0.1 (AntMaze, HumanoidMaze, Kitchen, VisualAntmaze, AntSoccer)
0.2 (Cube, Scene)
@@
In the online phase, our high-level policy is a Soft-Actor-Critic (SAC) agent (Haarnoja et al., 2018) with 10 critic networks, entropy backup disabled and LayerNorm added to the critics following the architecture design used in RLPD (Ball et al., 2023). We follow a similar strategy in ExPLORe (Li et al., 2024) where we sample 128 offline samples and 128 online samples and add RND reward bonus to all of the samples. The main difference is in the original ExPLORe paper is that they only add reward bonus to the offline data as additionally adding the bonus to the online replay buffer does not help for the maze goals they tested. In our experiments, we add the reward bonus to both offline data and online data, as it leads to better performance in goals where there is limited offline data coverage (see Appendix H.2).
@@ Appendix C Implementation details for baselines
Diffusion BC + JSRL. We use the diffusion model implementation from (Hansen-Estruch et al., 2023). Following the paper’s implementation, we train the model for 3 million gradient steps with a dropout rate of 0.1 and a cosine decaying learning rate schedule from the learning rate of $0.0003$. In the online phase, in the beginning of every episode, with probability $p$, we rollout the diffusion policy for a random number of steps that follows a geometric distribution $\mathrm{Geom}(1-\gamma)$ before sampling actions from the online agent (inspired by (Li et al., 2023)). A RND bonus is also added to the online batch on the fly with a coefficient of $2.0$ to further encourage online exploration of the SAC agent. The same coefficient is used in all other non-skill based baselines. For skill based baselines, we scale up the RND coefficient by the horizon length (4) to account for different reward scale. Following the BC + JSRL baseline used in ExPLORe (Li et al., 2024), we use a SAC agent with an ensemble of 10 critic networks, one actor network, with no entropy backup and LayerNorm in the critic networks. This configuration is used for all baselines on all environments. On AntMaze, We perform a hyperparameter sweep on both $p=\{0.5,0.75,0.9\}$ and the geometric distribution parameter $\gamma=\{0.99,0.995,0.997\}$ on the large maze with the top right goal and find that $p=0.9$ and $\gamma=0.99$ works the best. We also use these parameters for the Visual AntMaze experiments. On AntSoccer, we use the same $p$ but raise the discount $\gamma$ to match the environment discount rate of $0.995$. On HumanoidMaze, we perform a sweep over $p=\{0.5,0.75,0.9\}$ on humanoidmaze-medium-navigate-v0 and find that $p=0.75$ works best. We still use $\gamma=0.995$. On Scene and Cube, we do a similar sweep for $p$ on the first task of Scene, Single Cube, and Double Cube, and find that $p=0.5$ works best. We still use $\gamma=0.995$. For Kitchen, we perform a sweep on the parameter $p=\{0.2,0.5,0.75,0.9\}$ and find that 0.75 works best. We use $\gamma=0.99$. We take the minimum of one random critic for AntMaze, Visual AntMaze, AntSoccer, and HumanoidMaze, and the minimum of two random critics for Kitchen, Scene, and Cube. For all methods, we use the same image encoder used in RLPD (Ball et al., 2023) for the Visual AntMaze task, with a latent dimension of 50 (encoded image is a 50 dimensional vector), which is then concatenated with proprioceptive state observations.
ExPLORe. We directly use the open-source implementation from https://github.com/facebookresearch/ExPLORe/. The only difference we make is to adjust the RND coefficient from $1.0$ to $2.0$ and additionally add such bonus to the online replay buffer (the original method only adds to the offline data). Empirically, we find a slightly higher RND coefficient improves performance slightly. The SAC configuration is the same as that of the Diffusion BC + JSRL agent. We found that taking the minimum of one random critic on Visual AntMaze worked better for ExPLORe than taking the minimum of two, so all non-skill based baselines use this hyperparameter value.
Online RL with trajectory skills. This baseline is essentially our method but without using the trajectory encoder in the VAE to label trajectory segments (with high-level skill action labels), so all stated implementation decisions also apply to Ours. Instead, we treat it directly as a high-level RL problem with the low-level skill policy completely frozen. The SAC agent is the same as the previous agents, except for on Visual AntMaze, where taking the minimum of 2 critics from the ensemble leads to better performance for Ours, so we use this parameter setting for all skill-based benchmarks. We compute the high-level reward as the discounted sum of the rewards received every $H$ environment steps. During the $5\times 10^{3}$ steps before the start of training, we sample random actions from the state-based prior. For Visual AntMaze, we use the learned image encoder from the VAE to initialize both the critic image encoder and the RND network. If using ICVF, we initialize the RND network with the ICVF encoder instead.
Online RL with HILP skills. This baseline is the same as the one above but with the skills from a recent unsupervised offline skill discovery method, HILP (Park et al., 2024b). We use the official open-source implementation https://github.com/seohongpark/HILP and run the pretraining to obtain the skill policies. Then, we freeze the skill policies and learn a high-level RL agent to select skills every $H$ steps.
HILP w/ offline data. This novel baseline is the same as Online w/ HILP Skills, except that we also relabel the offline trajectories and use them as additional data for learning the high-level policy online (similar to our proposed method). To relabel trajectories with the estimated HILP skill, we compute the difference in the latent representation of the final state $s_{H}$ and initial state $s_{0}$ in the trajectory, so $\hat{z}\leftarrow\frac{\phi_{\mathrm{HILP}}(s_{H})-\phi_{\mathrm{HILP}}(s_{0})}{\|\phi_{\mathrm{HILP}}(s_{H})-\phi_{\mathrm{HILP}}(s_{0})\|_{2}}$. We normalize the skill vector since the pretrained HILP policies use a normalized vector as input. The high-level RL agent is the same as our method, except the skill relabeling is done using the latent difference rather than the trajectory encoder.
@@
Ours. We follow Li et al. (2024) (ExPLORe) to relabel offline data with optimistic reward estimates using RND and a reward model. For completeness, we describe the details below. We initialize two networks $g_{\phi}(s,z)$, $\bar{g}(s,z)$ that each outputs an $L$-dimensional feature vector predicted from the state and (tanh-squashed) high-level action. During online learning, $\bar{g}(s,z)$ is fixed and we only update the parameters of the other network $g_{\phi}(s,z)$ to minimize the $L_{2}$ distance between the feature vectors predicted by the two networks on the new high-level transition $(s^{\mathrm{new}}_{0},z^{\mathrm{new}},r^{\mathrm{new}},s^{\mathrm{new}}_{H})$:
@@
In addition to the two networks, we also learn a reward model $r_{\psi}(s,z)$ that minimizes the reward loss below on the transitions $(s_{0},z,r,s_{H})$ from online replay buffer:
$\mathcal{L}(\psi)=\|r_{\psi}(s_{0},z)-r\|_{2}^{2}.$
@@
$r_{\mathrm{UCB}}(s,z)\leftarrow r_{\psi}(s_{0},z)+\alpha\|g_{\phi}(s_{0},z)-\bar{g}(s_{0},z)\|_{2}^{2},$
@@
Figure 6: Success rate on Visual AntMaze environment with and without ICVF. Ours works well without ICVF, almost matching the original performance. However, the other baselines Online w/ Trajectory Skills and ExPLORe achieve far worse performance without ICVF, which shows that using offline data both for extracting skills and online learning leads to better utilization of noisy exploration bonuses. Initializing ExPLORe critic with ICVF helps, but does not substantially change performance.
Appendix D Domain Details

### v4

@@ 3 Problem Formulation
[…]
4 Skills from Unlabeled Prior Data for Exploration (SUPE)
In this section, we describe how we utilize the unlabeled trajectory dataset to accelerate online exploration. Our method, SUPE, can be divided into two parts. The first part is the offline pretraining phase where we extract skills from the offline data with a trajectory-segment VAE. The second part is the online learning phase where we train a high-level off-policy agent to compose the pretrained skills by leveraging examples from both the offline data and online replay buffer. Algorithm 1 describes our method.
Pretraining with trajectory VAE. Without knowledge of the downstream task, we need to capture the diverse range of multitask behavior from the unlabeled offline dataset as accurately as possible. At the same time, we aim to leverage this dataset to more efficiently learn a high-level policy which composes these skills to solve a particular unknown task online. We achieve this by adopting a trajectory VAE design from prior methods (Ajay et al., 2021; Pertsch et al., 2021) where a short trajectory segment $\tau_{[H]}=\{s_{0},a_{0},s_{1},\cdots,s_{H-1},a_{H-1}\}$ is first fed into a trajectory encoder $f_{\theta}(z|\tau)$ that outputs a distribution over the skill latent $z$. Then, a skill policy $\pi_{\theta}(a|s,z)$ is trained to reconstruct the actions in the segment using a latent sampled from the distribution outputted by the encoder. This design allows us to use the trajectory encoder $f_{\theta}(z|\tau)$ to obtain an estimate of the corresponding skill latent $\hat{z}$ for any trajectory segment. Following prior work (Pertsch et al., 2021), we learn a state-dependent prior $p_{\theta}(z|s)$ to help accommodate the difference in behavior diversity of different states. Putting them all together, the loss function is
$\mathcal{L}_{\theta}(\tau)$ $=\beta D_{\mathrm{KL}}(f_{\theta}(z|\tau)||p_{\theta}(z|s_{0}))-$ (1)
$\mathbb{E}_{z\sim f_{\theta}(z|\tau)}\left[\sum_{h=0}^{H-1}\log\pi_{\theta}(a_{h}|s_{h},z)\right].$
Online exploration with trajectory skills. To effectively leverage the pretrained trajectory skills, we learn a high-level off-policy agent that picks which skills to use to explore and solve the task efficiently. Following prior work on fixed-horizon skills (Pertsch et al., 2021), our high-level policy selects a latent skill $z$ every $H$ environment steps. To transform the unlabeled offline data into high-level, task-relevant transitions, we need both the high-level action and the reward label corresponding to each length $H$ trajectory segment $\tau_{[H]}$. For the high-level action, we use the trajectory encoder $f_{\theta}(z|\tau)$ to obtain an estimate of the skill latent $\hat{z}$. For the reward, we maintain an upper-confidence bound (UCB) estimate of the reward value for each state and skill latent pair $(s,z)$ inspired by prior work (Li et al., 2024) (where it does so directly in the state-action space $(s,a)$), and pseudo-label the transition with such an optimistic reward estimate. The reward estimate is recomputed before updating the high-level agent, since the estimate changes over time, while the trajectory encoding is computed before starting online learning, since this label does not change. The pseudo-labeling is summarized below:
Practical implementation details. Following prior work on trajectory-segment VAEs (Ajay et al., 2021; Pertsch et al., 2021), we use a Gaussian distribution (where both the mean and diagonal covariance are learnable) for the trajectory encoder, the skill policy, as well as the state-dependent prior. While Pertsch et al. (2021) use a KL constraint between the high-level policy and the state-dependent prior, we use a simpler design without the KL constraint that performs better in practice (as we show in Appendix G). To achieve this, we adapt the policy parameterization from Haarnoja et al. (2018), where the action value is enforced to be between $-1$ and $1$ using a $\mathrm{tanh}$ transformation, and entropy regularization is applied on the squashed space. We use this policy parameterization for the high-level policy $\pi(z|s)$ to predict the skill action in the squashed space $z_{\mathrm{sqaushed}}$. We then recover the actual skill action vector by unsquashing it according to $z=\mathrm{arctanh}(z_{\mathrm{sqaushed}})$, so that it can be used by our skill policy $\pi_{\theta}(a|s,z)$. For upper-confidence bound (optimistic) estimation of the reward, ($r_{\mathrm{UCB}}(s_{0},\hat{z})$), we directly borrow the UCB estimation implementation in Li et al. (2024) (Section 3, practical implementation section in their paper), where they use a combination of the random network distillation (RND) (Burda et al., 2018) reward bonus and the predicted reward from a reward model (see Appendix C for more details). For the off-policy high-level agent, we follow Li et al. (2024) to use RLPD (Ball et al., 2023) that takes a balanced number of samples from the offline data and the online replay buffer for agent optimization. In addition to using the optimistic offline reward label, we also add the RND bonus to the online batch as we have found it to further accelerate online learning.
(b) b) humanoidmaze-medium/large/giant.
(c) c) antsoccer-arena/medium
(d) d) kitchen
(e) e) cube
(f) f) scene
(g) g) visual-antmaze
Figure 2: We experiment on 8 challenging, sparse-reward domains. a): antmaze with three different layouts and the corresponding four goal locations (denoted as the red dots); b): humanoidmaze with three layouts; c): antsoccer with two layouts d): kitchen; e): cube-single and cube-double; f): scene; g): visual-antmaze with colored floor and local $64\times 64$ image observations.
5 Experimental Results
We present a series of experiments to evaluate the effectiveness of our method in discovering fast exploration strategies. We specifically focus on long-horizon, sparse-reward settings, where online exploration is especially important. In particular, we aim to answer the following questions:
  1. 1.
    Can SUPE leverage pretrained skills to accelerate online learning?
  2. 2.
    Can SUPE find goals faster than prior methods?
  3. 3.
    How sensitive is SUPE to hyperparameters?
5.1 Experimental setup
We conduct our experiments on 8 challenging sparse-reward domains and provide a brief description of each of the domains below (more details available in Appendix E).
State-based locomotion: antmaze, humanoidmaze, antsoccer. The first set of domains involve controlling and navigating a robotic agent in a complex environment. antmaze is a standard benchmark for offline-to-online RL from D4RL (Fu et al., 2020). humanoidmaze and antsoccer are locomotion domains from OGBench, a offline goal-conditioned RL benchmark (Park et al., 2024a). In antmaze and humanoid, a humanoid or ant agent must learn to navigate to various goal locations. antsoccer adds the difficulty of moving a soccer ball to a desired position. Each of the antmaze and humanoidmaze domains has three different maze layouts. For antmaze, we test on four different goal locations for each maze layout. For humanoidmaze and antsoccer we test on one goal. For humanoidmaze, we use both the navigate and stitch datasets. For antsoccer, we use the navigate dataset for the arena and medium maze layouts.
State-based manipulation: kitchen, cube-single, cube-double, scene. Next, we consider a set of manipulation domains that require a wide range of manipulation skills. kitchen is a standard benchmark from D4RL where a robotic arm needs to complete a set of manipulation tasks (e.g., turn on the microwave, move the kettle) in sequence in a kitchen scene. cube-single, cube-double, and scene are three domains from OGBench (Park et al., 2024a). For both cube-* domains, the robotic arm must arrange one or more cube objects to desired goal locations (e.g., stacking on top of each other) in tasks that mainly involve composing pick and place motions. For scene, the robotic arm can interact with a more diverse set of objects: a window, a drawer, a cube and two locks that control the window and the drawer. The tasks in scene are also longer, requiring the composition of multiple atomic behaviors (e.g., locking and unlocking, opening the drawer/window, moving the cube).
Pixel-based: visual-antmaze is a challenging visual domain introduced by Park et al. (2023a). The agent has all the same proprioceptive information from the state-based antmaze environment except the agent’s $(x,y)$ position, which the agent must obtain from a $64\times 64$ image observation of its surroundings.
Figure 3: Aggregated normalized return across eight different domains. We evaluate baselines that use offline data only for pretraining (Refer to caption), only during online learning (Refer to caption), and our methods that use offline data for both pretraining and online learning (Refer to caption). SUPE matches or outperforms all other methods on all domains except antsoccer and scene, where SUPE (HILP) performs better. Section 5.2 contains more details about the baselines. For kitchen, we use 16 seeds. For all other domains, we use 8 seeds.
To evaluate our method on these domains, we take the datasets in these benchmarks and remove the reward label as well as any information in the transition that may reveal the information about the termination of an episode. For all of the domains above, we use the normalized return, a standard metric for D4RL (Fu et al., 2020) environments, as the main evaluation metric. For the kitchen domain, the normalized return represents the average percentage of the tasks that are solved. For tasks in other domains, the normalized return represents the average task success rate. For all our plots, the shaded area indicates the 95% stratified bootstrap confidence interval and the solid line indicates the interquartile mean following Agarwal et al. (2021). See more details about the tasks and rewards in Appendix E.
Figure 4: Interquartile mean (IQM) of the number of environment steps taken to reach the goal (lower the better). The first goal time is considered to be the maximum ($0.3\times 10^{6}$ steps) if the agent never finds the goal. SUPE is the most consistent, achieving performance better than all other baselines on all layouts (4 tasks/goals for each layout and 8 seeds for each task).
5.2 Comparisons
While there are no existing methods that pretrain on unlabeled offline trajectories and use these trajectories for online learning, there are methods that use offline data for either purpose. We first consider a baseline that directly performs online learning with no pretraining.
ExPLORe (Li et al., 2024). Unlike our method, this baseline does not perform unsupervised skill pretraining and directly learns a 1-step policy online. As we will show, pretraining is crucial for success on more challenging domains. Like our method, ExPLORe uses an exploration bonus on offline data to encourage exploration. It is worth noting that the original ExPLORe method does not make use of an online reward bonus and only uses the RND network to optimistically label offline data. We additionally add the reward bonus to the online batch from the replay buffer to help the agent explore beyond the offline data distribution. For completeness, we include the performance of the original ExPLORe in Figure 12. For all the baselines below (including our method), we add the RND bonus to the online training batch to ensure a fair comparison.
We then consider baselines that pretrain on unlabeled offline data but do not use the data online.
DBC+JSRL. We pretrain an expressive diffusion policy to imitate the offline data via a behavior cloning (BC) loss. At the beginning of each online episode, we rollout this pretrained policy for a random number of steps from the initial state before using switching to the online RL agent (Uchendu et al., 2023; Li et al., 2023). One might expect that a sufficiently expressive policy class can capture the action distribution of the offline data accurately enough to provide an informative initial state distribution for online exploration. This baseline is an upgraded version of the BC+JSRL baseline used in ExPLORe (Li et al., 2024), with the Gaussian BC policy replaced by a diffusion policy.
Trajectory Skills and HILP Skills. We also consider two baselines that pretrain skills offline, but discard the offline data during online learning, training the high-level policy from scratch. Each baseline uses a different skill pretraining method. In addition to the trajectory skill approach used by our method, we also consider a recently proposed unsupervised offline skill discovery method where skills are pretrained to traverse a learned Hilbert representation space (Park et al., 2024b) (HILP). For both baselines, we use the same high-level RL agent as our method, but learn this high-level policy purely from online interaction with the environment. While learning a high-level policy from scratch to compose trajectory VAE skills is similar to SPiRL (Pertsch et al., 2021), our baseline features two additional improvements. The first improvement is replacing the KL constraint with entropy regularization (same as SUPE as described in Section 4, practical implementation details). The second improvement is the online RND bonus that is also added to all other methods.
Finally, we introduce a novel baseline that also uses offline data during pretraining and online exploration, but uses HILP skills rather than trajectory-based skills.
SUPE (HILP). Like trajectory skills, we can also estimate the corresponding HILP skill latent for any trajectory segment. HILP learns a latent space of the observations (via an encoder $\phi_{\mathrm{HILP}}$) and learns skills that move the agent in a certain direction $z$ in the latent space. For any high-level transition $(s_{0},s_{H})$, we can simply take $\hat{z}\leftarrow\mathrm{normalize}(\phi_{\mathrm{HILP}}(s_{H})-\phi_{\mathrm{HILP}}(s_{0}))$, the normalized difference vector that points from $s_{0}$ to $s_{H}$ in the latent space. Aside from skill labeling, we use the exact same high-level RL agent as used in our method.
For visual-antmaze, we use the same image encoder used in RLPD (Ball et al., 2023). We also follow one of our baselines, ExPLORe (Li et al., 2024), to use ICVF (Ghosh et al., 2023), a method that uses task-agnostic value functions to learn image/state representations from passive data. ICVF takes in an offline unlabeled trajectory dataset with image observations and pretrain an image encoder in an unsupervised manner. Following ExPLORe, we take the weights of the image encoder from ICVF pretraining to initialize the image encoder’s weights in the RND network. To make the comparison fair, we also apply ICVF to all baselines (details in Appendix F).
Figure 5: SUPE skill policy rollouts in humanoid-maze-(medium/large/giant). For each trajectory, we first sample a skill latent, and then repeatedly sample a lower-level action from the low-level skill policy conditioned on that fixed latent. We plot the $x$-$y$ position of the agent throughout each of 16 such trajectories for each maze.
Figure 6: Sensitivity analysis on the RND coefficient ($\alpha$) and the skill horizon length ($H$) on a subset of tasks. For different hyperparameter values, we report the interquartile mean (IQM) of the normalized return across seven representative tasks. The performance of SUPE is not very sensitive to the magnitude of $\alpha$ as long as it is within a reasonable range $(2,16)$. Without the bonus ($\alpha=0$), SUPE performs significantly worse. A skill horizon length of 4 performs significantly better than a horizon length of 2 or 8. We use a skill length of 4 and $\alpha=8$ for all experiments with skill-based methods. The normalized return for each individual task we use for this analysis can be found in Appendix H. We average over 4 seeds.
5.3 Can SUPE leverage pretrained skills to accelerate online learning?
Figure 3 shows the aggregated performance of our approach on all eight domains. Our method outperforms all prior methods on all domains except antsoccer, where HILP Skills performs better, and cube-single, where ExPLORe performs competitively. Additionally, we note that while SUPE (HILP) performs best on scene and antsoccer, SUPE performs best on all other tasks. This suggests that different skill pretraining methods may excel at different tasks. Unlike all prior methods, both SUPE and SUPE (HILP) leverage offline data twice (both during offline pretraining and online learning), which demonstrates how this new approach is critical to surpassing prior methods. Both skill-based online learning baselines (Trajectory Skills and HILP Skills) each consistently perform worse than the corresponding novel method that uses offline data during online learning (SUPE and SUPE (HILP)), which further demonstrates the importance of using offline data twice. ExPLORe uses offline data during online learning, but does not pretrain skills, leading to slower learning on all 8 domains and difficulty achieving any significant return on any domains other than the easier antmaze and cube-single tasks. We also report the performance on individual antmaze mazes and kitchen tasks in Figure 11, and observe that our method outperforms the baselines more on harder environments. This trend continues with humanoidmaze, where SUPE is the only method to achieve nonzero final return on the more difficult large and giant mazes (see Appendix J.1, Figure 16). These experiments suggest that pretraining skills from prior data and leveraging offline data in the online phase are both crucial for sample-efficient learning, especially in more challenging environments.
To provide more insight on the sample-efficiency of our method, we also provide a visualization of rollouts from the skill policy in Figure 5. We notice that some of the frozen skill policies are able to navigate quite far from the starting point, which suggests that such structured navigation behaviors allow the high-level policy to effectively explore with a reduced time horizon.
For the visual domain, we perform an ablation study to assess the importance of the ICVF pretrained representation, which we include in Appendix F. While ICVF combines synergistically with our method to further accelerate learning, initializing RND image encoder weights using ICVF is not critical to its success.
5.4 Can SUPE find goals faster than prior methods?
While we have demonstrated that our method outperforms prior methods, it is still unclear if our method actually leads to better exploration (instead of simply learning the high-level policy better). In this section, we study the exploration aspect in isolation in the antmaze domain. Figure 4 reports the number of online environment interaction steps for the agent to reach the goal for the first time. Such a metric allows us to assess how efficiently the agent explores in the maze. Figure 4 shows that our method reaches goals faster than every baseline on all three maze layouts, which confirms that our method not only learns faster, but does so by exploring more efficiently. We include the results for each individual goal location and maze layout in the Appendix (Table 3) for completeness.
5.5 How sensitive is SUPE to hyperparameters?
Finally, we analyze the sensitivity of our method to hyperparameter selection. We focus on two hyperparameters, 1) $\alpha$: the amount of optimism (RND coefficient) used when labeling offline data with UCB rewards (Equation 2), and 2) $H$: the length of the skill. We select one representative task from each state-based domain, and study how different $\alpha$ and $H$ values affect the performance on these tasks (Figure 6). When the RND bonus is removed ($\alpha=0$), our method performs significantly worse, highlighting the importance of optimistic labeling on efficient online exploration. While we observe some variability across individual tasks (Appendix H, Figure 9), our method is largely not very sensitive to the RND coefficient value. The aggregated performance is similar for $\alpha\in\{2,8,16\}$. Another key hyperparameter in our method is the skill horizon length (see Figure 6, right). We find that while there is some variability across individual tasks (Appendix H, Figure 10), a skill horizon length of 4 generally performs the best, and shorter or longer horizons perform much worse.
6 Discussion and Limitations
In this work, we propose SUPE, a method that leverages unlabeled offline trajectory data to accelerate online exploration and learning. The key insight is to use unlabeled trajectories twice, to 1) extract a set of low-level skills offline, and 2) serve as additional data for a high-level off-policy RL agent to compose these skills to explore in the environment. This allows us to effectively combine the strengths from unsupervised skill pretraining and sample-efficient online RL methods to solve a series of challenging long-horizon sparse reward tasks significantly more efficiently than existing methods. Our work opens up avenues in making full use of prior data for scalable, online RL algorithms. First, our pretrained skills remain frozen during online learning, which may hinder online learning when the skills are not learned well or need to be updated as the learning progresses (see analysis in Appendix K and a failure case in Figure 19). Such problems could be alleviated by utilizing a better skill pretraining method, or allowing the low-level skills to be fine-tuned online. In addition, our approach relies on RND to maintain an upper confidence bound on the optimistic reward estimate. Although we find that RND works without ICVF on high-dimensional image observations in visual-antmaze, the use of RND in other high dimensional environments may require more careful consideration. Possible future directions include examining alternative methods of maintaining this bound.
Acknowledgments
Impact Statement
This paper presents work whose goal is to advance the field of reinforcement learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Compute Resources
We ran all our experiments on NVIDIA A5000 GPU and V100 GPUs. In total, the results in this paper required approximately $16600$ GPU hours.
@@ Appendix B VAE Architecture and hyperparameters
GRU Hidden Size 256 (antmaze, kitchen, visual-antmaze)
$512$ (antsoccer, humanoidmaze, scene, cube-*)
GRU Layers $2$ hidden layers (antmaze, kitchen, visual-antmaze)
$3$ hidden layers (antsoccer, humanoidmaze, scene, cube-*)
KL Coefficient ($\beta)$ $0.1$ (antmaze, humanoidmaze, kitchen, visual-antmaze, antsoccer)
$0.2$ (cube-*, scene)
@@
In the online phase, our high-level policy is a Soft-Actor-Critic (SAC) agent (Haarnoja et al., 2018) with 10 critic networks, entropy backup disabled and LayerNorm added to the critics following the architecture design used in RLPD (Ball et al., 2023). Similar to ExPLORe (Li et al., 2024), we sample 128 offline samples and 128 online samples and add an RND reward bonus to all of the samples. While the original ExPLORe paper only added a reward bonus to the offline data, we add the reward bonus to both the offline data and online data, as it leads to better performance in goals where there is limited offline data coverage (see Appendix I.2).
@@ Appendix C Implementation details for baselines
SUPE. We follow Li et al. (2024) (ExPLORe) to relabel offline data with optimistic reward estimates using RND and a reward model. For completeness, we describe the details below. We initialize two networks $g_{\phi}(s,z)$, $\bar{g}(s,z)$ that each outputs an $256$-dimensional feature vector predicted from the state and (tanh-squashed) high-level action. During online learning, $\bar{g}(s,z)$ is fixed and we only update the parameters of the other network $g_{\phi}(s,z)$ to minimize the $L_{2}$ distance between the feature vectors predicted by the two networks on the new high-level transition $(s^{\mathrm{new}}_{0},z^{\mathrm{new}},r^{\mathrm{new}},s^{\mathrm{new}}_{H})$:
@@
In addition to the two networks, we also learn a reward model $r_{\zeta}(s,z)$ that minimizes the reward loss below on the transitions $(s_{0},z,r,s_{H})$ from online replay buffer:
$\mathcal{L}(\zeta)=\|r_{\zeta}(s_{0},z)-r\|_{2}^{2}.$
@@
$r_{\mathrm{UCB}}(s,z)\leftarrow r_{\zeta}(s_{0},z)+\alpha\|g_{\phi}(s_{0},z)-\bar{g}(s_{0},z)\|_{2}^{2},$
@@
For online training batch from the replay buffer, we also add online reward bonus provided by RND with the same $\alpha$ coefficient. For all the baselines below, we also add the online RND bonus to the online batch to encourage exploration.
Online Only. In the appendix, we additionally include a baseline that does not use offline data at all. This baseline directly runs the SAC agent from scratch online with no pretraining and no usage of the offline data.
DBC+JSRL. We use the diffusion model implementation from (Hansen-Estruch et al., 2023). Following the paper’s implementation, we train the model for 3 million gradient steps with a dropout rate of 0.1 and a cosine decaying learning rate schedule from the learning rate of $0.0003$. In the online phase, in the beginning of every episode, with probability $p$, we rollout the diffusion policy for a random number of steps that follows a geometric distribution $\mathrm{Geom}(1-\gamma)$ before sampling actions from the online agent (inspired by (Li et al., 2023)). A RND bonus is also added to the online batch on the fly with a coefficient of $2.0$ to further encourage online exploration of the SAC agent. The same coefficient is used in all other non-skill based baselines. For skill based baselines, we scale up the RND coefficient by the horizon length (4) to account for different reward scale. Following the BC+JSRL baseline used in ExPLORe (Li et al., 2024), we use a SAC agent with an ensemble of 10 critic networks, one actor network, with no entropy backup and LayerNorm in the critic networks. This configuration is used for all baselines on all environments. On antmaze, we perform a hyperparameter sweep on both $p=\{0.5,0.75,0.9\}$ and the geometric distribution parameter $\gamma=\{0.99,0.995,0.997\}$ on the large maze with the top right goal and find that $p=0.9$ and $\gamma=0.99$ works the best. We also use these parameters for the Visual antmaze experiments. On antsoccer, we use the same $p$ but raise the discount $\gamma$ to match the environment discount rate of $0.995$. On humanoidmaze, we perform a sweep over $p=\{0.5,0.75,0.9\}$ on humanoidmaze-medium-navigate and find that $p=0.75$ works best. We still use $\gamma=0.995$. On Scene and Cube, we do a similar sweep for $p$ on the first task of Scene, Single Cube, and Double Cube, and find that $p=0.5$ works best. We still use $\gamma=0.995$. For kitchen, we perform a sweep on the parameter $p=\{0.2,0.5,0.75,0.9\}$ and find that 0.75 works best. We use $\gamma=0.99$. We take the minimum of one random critic for antmaze, visual-antmaze, antsoccer, and humanoidmaze, and the minimum of two random critics for kitchen, scene, and cube. For antsoccer, we performed a sweep on antsoccer-arena-navigate for each method over taking a minimum of 1 or 2 random critics, and found that taking a minimum of 1 critic was always better. For all methods, we use the same image encoder used in RLPD (Ball et al., 2023) for the visual-antmaze task, with a latent dimension of 50 (encoded image is a 50 dimensional vector), which is then concatenated with proprioceptive state observations.
ExPLORe. We directly use the official open-source implementation. The only difference we make is to adjust the RND coefficient from $1.0$ to $2.0$ and additionally add such bonus to the online replay buffer (the original method only adds to the offline data). Empirically, we find a slightly higher RND coefficient improves performance slightly. The SAC configuration is the same as that of the DBC+JSRL baseline. We found that taking the minimum of one random critic on visual-antmaze worked better for ExPLORe than taking the minimum of two, so all non-skill based baselines use this hyperparameter value.
Trajectory Skills. This baseline is essentially our method but without using the trajectory encoder in the VAE to label trajectory segments (with high-level skill action labels), so all stated implementation decisions also apply to SUPE. Instead, we treat it directly as a high-level RL problem with the low-level skill policy completely frozen. The SAC agent is the same as the previous agents, except for on Visual AntMaze, where taking the minimum of 2 critics from the ensemble leads to better performance for SUPE, so we use this parameter setting for all skill-based benchmarks. We compute the high-level reward as the discounted sum of the rewards received every $H$ environment steps. During the $5\times 10^{3}$ steps before the start of training, we sample random actions from the state-based prior. For Visual AntMaze, we use the learned image encoder from the VAE to initialize both the critic image encoder and the RND network. If using ICVF, we initialize the RND network and critic encoder with the ICVF encoder instead.
HILP Skills. This baseline is the same as the one above but with the skills from a recent unsupervised offline skill discovery method, HILP (Park et al., 2024b). We use the official open-source implementation from the authors and run the pretraining to obtain the skill policies. Then, we freeze the skill policies and learn a high-level RL agent to select skills every $H$ steps.
SUPE (HILP). This novel baseline is the same as HILP Skills, except that we also relabel the offline trajectories and use them as additional data for learning the high-level policy online (similar to our proposed method). To relabel trajectories with the estimated HILP skill, we compute the difference in the latent representation of the final state $s_{H}$ and initial state $s_{0}$ in the trajectory, so $\hat{z}\leftarrow\frac{\phi_{\mathrm{HILP}}(s_{H})-\phi_{\mathrm{HILP}}(s_{0})}{\|\phi_{\mathrm{HILP}}(s_{H})-\phi_{\mathrm{HILP}}(s_{0})\|_{2}}$. We normalize the skill vector since the pretrained HILP policies use a normalized vector as input. The high-level RL agent is the same as our method, except the skill relabeling is done using the latent difference rather than the trajectory encoder.
Appendix D Additional Related Work
Options framework. Many existing works on building hierarchical agents also adopt the options framework (Sutton et al., 1999; Menache et al., 2002; Chentanez et al., 2004; Mannor et al., 2004; Şimşek & Barto, 2004, 2007; Konidaris, 2011; Daniel et al., 2016; Srinivas et al., 2016; Fox et al., 2017; Bacon et al., 2017; Kim et al., 2019; Bagaria & Konidaris, 2019; Bagaria et al., 2024). The options framework provides a way to learn skills with varying time horizons, often defined by learnable initiation and/or termination conditions (Sutton et al., 1999). We use skills with a fixed horizon length because it allows us to avoid the additional complexity of learning initiation or termination conditions, and frame the skill pretraining phase as a simple supervised learning task.
Reinforcement learning with observation-only offline data. In this problem setting, the offline data only contains observations (e.g., videos) (Ma et al., 2022; Ghosh et al., 2023; Song et al., 2024). The closest to our work are Ma et al. (2022) and Ghosh et al. (2023), which focus on pre-training on observation-only offline data to extract good image representations that are suitable for down-stream RL tasks. Different from their work, we pre-train low-level skills rather than representations for online learning.
Appendix E Domain Details